Two parallel plates at a distance of 8.0*10^-5 apart are maintained at a p.d of 600 volts with the negative plate earthed.what is the electric field strength?
1
Expert's answer
2020-06-09T13:19:09-0400

We can find the electric field strength from the formula:


"E = \dfrac{V}{d} = \dfrac{600 \ V}{8.0 \cdot 10^{-5} \ m} = 7.5 \cdot 10^6 \ \dfrac{V}{m}."

Answer:

"E = 7.5 \cdot 10^6 \ \dfrac{V}{m}."